# Order-Optimal Global Convergence for Average Reward Reinforcement Learning via Actor-Critic Approach

### v2

@@ 5.4 Critic and Average Reward Analysis
$(a)~{}\left\lVert\operatorname{\mathbb{E}}_{k,h}\left[\hat{A}^{\mathrm{MLMC}}_{\mathbf{v},k,h}\right]-A_{\mathbf{v}}(\theta_{k})\right\rVert^{2}\leq\operatorname*{\mathcal{O}}\left(c_{\beta}^{2}t_{\mathrm{mix}}T_{\max}^{-1}\right)$
$(b)~{}\operatorname{\mathbb{E}}_{k,h}\left[\left\lVert\hat{A}^{\mathrm{MLMC}}_{\mathbf{v},k,h}-A_{\mathbf{v}}(\theta_{k})\right\rVert^{2}\right]\leq\operatorname*{\mathcal{O}}\left(c_{\beta}^{2}t_{\mathrm{mix}}\log T_{\max}\right)$
$(c)~{}\left\lVert\operatorname{\mathbb{E}}_{k,h}\left[\hat{b}^{\mathrm{MLMC}}_{\mathbf{v},k,h}\right]-b_{\mathbf{v}}(\theta_{k})\right\rVert^{2}\leq\operatorname*{\mathcal{O}}\left(c_{\beta}^{2}t_{\mathrm{mix}}T_{\max}^{-1}\right)$
$(d)~{}\operatorname{\mathbb{E}}_{k,h}\left[\left\lVert\hat{b}^{\mathrm{MLMC}}_{\mathbf{v},k,h}-b_{\mathbf{v}}(\theta_{k})\right\rVert^{2}\right]\leq\operatorname*{\mathcal{O}}\left(c_{\beta}^{2}t_{\mathrm{mix}}\log T_{\max}\right)$
@@
$\mathbb{E}_{k}\left[\|\xi_{k}-\xi_{k}^{*}\|^{2}\right]\leq\frac{1}{H^{2}}\|\xi^{k}_{0}-\xi_{k}^{*}\|^{2}+\tilde{\operatorname*{\mathcal{O}}}\left(\frac{c_{\beta}^{4}t_{\mathrm{mix}}}{\lambda^{4}H}\right),$
$\|\operatorname{\mathbb{E}}_{k}[\xi_{k}]-\xi_{k}^{*}\|^{2}\leq\operatorname*{\mathcal{O}}\left(\dfrac{c_{\beta}^{2}t_{\mathrm{mix}}}{\lambda^{2}H^{2}}\left\lVert\xi_{0}^{k}-\xi_{k}^{*}\right\rVert^{2}+\dfrac{c_{\beta}^{6}t^{2}_{\mathrm{mix}}}{\lambda^{6}H^{2}}\right)$
@@ Appendix A Proof of Lemma 1
The proof of the lemma is straightforward and follows a similar structure to the one in [Mondal and Aggarwal, 2024]. The major distinction is that the bound derived in [Mondal and Aggarwal, 2024] applies to the discounted reward setting, whereas our derivation pertains to the average-reward case. We begin by stating a useful lemma.
Lemma 4 (Lemma 4, [Bai et al., 2024]).
@@
$\operatorname{\mathbb{E}}_{s\sim d^{\pi^{*}}}[KL(\pi^{*}(\cdot|s)\|\pi_{\theta_{k}}(\cdot|s))-KL(\pi^{*}(\cdot|s)\|\pi_{\theta_{k+1}}(\cdot|s))]$
@@
$\begin{split}&\operatorname{\mathbb{E}}\left[\operatorname{\mathbb{E}}_{s\sim d^{\pi^{*}}}\left[KL(\pi^{*}(\cdot|s)\|\pi_{\theta_{k}}(\cdot|s))-KL(\pi^{*}(\cdot|s)\|\pi_{\theta_{k+1}}(\cdot|s))\right]\right]\\ &\geq\alpha[J^{*}-\operatorname{\mathbb{E}}\left[J(\theta_{k})]\right]-\alpha\sqrt{\epsilon_{\mathrm{bias}}}\\ &\hskip 56.9055pt+\alpha\operatorname{\mathbb{E}}\left[\operatorname{\mathbb{E}}_{s\sim d^{\pi^{*}}}\operatorname{\mathbb{E}}_{a\sim\pi^{*}(\cdot|s)}[\nabla_{\theta}\log\pi_{\theta_{k}}(a|s)\cdot(\operatorname{\mathbb{E}}_{k}[\omega_{k}]-\omega^{*}_{k})]\right]-\frac{G_{2}\alpha^{2}}{2}\operatorname{\mathbb{E}}\left[\|\omega_{k}\|^{2}\right]\\ &\overset{}{\geq}\alpha[J^{*}-\operatorname{\mathbb{E}}\left[J(\theta_{k})]\right]-\alpha\sqrt{\epsilon_{\mathrm{bias}}}\\ &\hskip 56.9055pt-\alpha\operatorname{\mathbb{E}}\left[\operatorname{\mathbb{E}}_{s\sim d^{\pi^{*}}}\operatorname{\mathbb{E}}_{a\sim\pi^{*}(\cdot|s)}[\|\nabla_{\theta}\log\pi_{\theta_{k}}(a|s)\|\|\operatorname{\mathbb{E}}_{k}[\omega_{k}]-\omega^{*}_{k}\|]\right]-\frac{G_{2}\alpha^{2}}{2}\operatorname{\mathbb{E}}\left[\|\omega_{k}\|^{2}\right]\\ &\overset{(a)}{\geq}\alpha[J^{*}-\operatorname{\mathbb{E}}\left[J(\theta_{k})]\right]-\alpha\sqrt{\epsilon_{\mathrm{bias}}}-\alpha G_{1}\operatorname{\mathbb{E}}\|(\operatorname{\mathbb{E}}_{k}[\omega_{k}]-\omega^{*}_{k})\|-\frac{G_{2}\alpha^{2}}{2}\operatorname{\mathbb{E}}\left[\|\omega_{k}\|^{2}\right]\end{split}$ (36)
@@
$\begin{split}J^{*}-\operatorname{\mathbb{E}}[J(\theta_{k})]&\leq\sqrt{\epsilon_{\mathrm{bias}}}+G_{1}\operatorname{\mathbb{E}}\|(\operatorname{\mathbb{E}}_{k}[\omega_{k}]-\omega^{*}_{k})\|+\frac{G_{2}\alpha}{2}\operatorname{\mathbb{E}}\|\omega_{k}\|^{2}\\ &+\frac{1}{\alpha}\operatorname{\mathbb{E}}\left[\operatorname{\mathbb{E}}_{s\sim d^{\pi^{*}}}[KL(\pi^{*}(\cdot|s)\|\pi_{\theta_{k}}(\cdot|s))-KL(\pi^{*}(\cdot|s)\|\pi_{\theta_{k+1}}(\cdot|s))]\right]\end{split}$ (37)
@@ Appendix C Properties of the MLMC Estimates
This section provides some guarantees on the error of the MLMC estimator. This is similar to the results available in [Dorfman and Levy, 2022, Suttle et al., 2023, Beznosikov et al., 2023], although [Dorfman and Levy, 2022, Beznosikov et al., 2023] consider the case of unbiased estimates while our results deal with biased estimates.
@@
Lemma 6 (Lemma 1, [Beznosikov et al., 2023]).
@@ Appendix F Proof of Theorem 3
[Wei et al., 2020, Lemma 14] For any ergodic MDP with mixing time $t_{\mathrm{mix}}$, the following holds for any policy $\pi$.
@@ Appendix H Proof of Theorem 1
$\begin{split}J^{*}-\frac{1}{K}\sum_{k=0}^{K-1}&\operatorname{\mathbb{E}}[J(\theta_{k})]\leq\sqrt{\epsilon_{\mathrm{bias}}}+\frac{G_{1}}{K}\sum_{k=0}^{K-1}\operatorname{\mathbb{E}}\|(\operatorname{\mathbb{E}}_{k}\left[\omega_{k}\right]-\omega^{*}_{k})\|+\dfrac{\alpha G_{2}}{K}\sum_{k=0}^{K-1}\operatorname{\mathbb{E}}\|\omega_{k}-\omega_{k}^{*}\|^{2}\\ &+\dfrac{\alpha\mu^{-2}}{K}\sum_{k=0}^{K-1}\operatorname{\mathbb{E}}\|\nabla_{\theta}J(\theta_{k})\|^{2}+\frac{1}{\alpha K}\operatorname{\mathbb{E}}_{s\sim d^{\pi^{*}}}[KL(\pi^{*}(\cdot|s)\|\pi_{\theta_{0}}(\cdot|s))].\end{split}$ (57)
@@
We get the desired result by substituting the values of $H,K$ as stated in the theorem. We want to emphasize that the $G_{1}^{2}$ factor with the $\sqrt{\epsilon_{\mathrm{app}}}$ term is a standard component in actor-critic results with a linear critic [Suttle et al., 2023, Patel et al., 2024]. However, this factor is often not explicitly mentioned in previous works, whereas we have included it here.

### v3

@@ 5.4 Critic and Average Reward Analysis
[…]
@@
[…]
@@ 6 Conclusions
Impact Statement
The goal of this paper is to advance the current understanding of the field of Machine Learning. We do not see any potential societal consequences of our work.
@@ Appendix A Proof of Lemma 1
The proof of the lemma is inspired by the analysis in (Mondal & Aggarwal, 2024). The major distinction is that the bound derived in (Mondal & Aggarwal, 2024) applies to the discounted reward setting, whereas our derivation pertains to the average-reward case. We begin by stating a useful lemma.
Lemma 4 (Lemma 4, (Bai et al., 2024)).
@@
$\operatorname{\mathbb{E}}_{s\sim d^{\pi^{*}}}[\mathrm{KL}(\pi^{*}(\cdot|s)\|\pi_{\theta_{k}}(\cdot|s))-\mathrm{KL}(\pi^{*}(\cdot|s)\|\pi_{\theta_{k+1}}(\cdot|s))]$
@@
$\begin{split}&\operatorname{\mathbb{E}}\left[\operatorname{\mathbb{E}}_{s\sim d^{\pi^{*}}}\left[\mathrm{KL}(\pi^{*}(\cdot|s)\|\pi_{\theta_{k}}(\cdot|s))-\mathrm{KL}(\pi^{*}(\cdot|s)\|\pi_{\theta_{k+1}}(\cdot|s))\right]\right]\\ &\geq\alpha[J^{*}-\operatorname{\mathbb{E}}\left[J(\theta_{k})]\right]-\alpha\sqrt{\epsilon_{\mathrm{bias}}}\\ &\hskip 56.9055pt+\alpha\operatorname{\mathbb{E}}\left[\operatorname{\mathbb{E}}_{s\sim d^{\pi^{*}}}\operatorname{\mathbb{E}}_{a\sim\pi^{*}(\cdot|s)}[\nabla_{\theta}\log\pi_{\theta_{k}}(a|s)\cdot(\operatorname{\mathbb{E}}_{k}[\omega_{k}]-\omega^{*}_{k})]\right]-\frac{G_{2}\alpha^{2}}{2}\operatorname{\mathbb{E}}\left[\|\omega_{k}\|^{2}\right]\\ &\overset{}{\geq}\alpha[J^{*}-\operatorname{\mathbb{E}}\left[J(\theta_{k})]\right]-\alpha\sqrt{\epsilon_{\mathrm{bias}}}\\ &\hskip 56.9055pt-\alpha\operatorname{\mathbb{E}}\left[\operatorname{\mathbb{E}}_{s\sim d^{\pi^{*}}}\operatorname{\mathbb{E}}_{a\sim\pi^{*}(\cdot|s)}[\|\nabla_{\theta}\log\pi_{\theta_{k}}(a|s)\|\|\operatorname{\mathbb{E}}_{k}[\omega_{k}]-\omega^{*}_{k}\|]\right]-\frac{G_{2}\alpha^{2}}{2}\operatorname{\mathbb{E}}\left[\|\omega_{k}\|^{2}\right]\\ &\overset{(a)}{\geq}\alpha[J^{*}-\operatorname{\mathbb{E}}\left[J(\theta_{k})]\right]-\alpha\sqrt{\epsilon_{\mathrm{bias}}}-\alpha G_{1}\operatorname{\mathbb{E}}\|(\operatorname{\mathbb{E}}_{k}[\omega_{k}]-\omega^{*}_{k})\|-\frac{G_{2}\alpha^{2}}{2}\operatorname{\mathbb{E}}\left[\|\omega_{k}\|^{2}\right]\end{split}$ (37)
@@
$\begin{split}J^{*}-\operatorname{\mathbb{E}}[J(\theta_{k})]&\leq\sqrt{\epsilon_{\mathrm{bias}}}+G_{1}\operatorname{\mathbb{E}}\|(\operatorname{\mathbb{E}}_{k}[\omega_{k}]-\omega^{*}_{k})\|+\frac{G_{2}\alpha}{2}\operatorname{\mathbb{E}}\|\omega_{k}\|^{2}\\ &+\frac{1}{\alpha}\operatorname{\mathbb{E}}\left[\operatorname{\mathbb{E}}_{s\sim d^{\pi^{*}}}[\mathrm{KL}(\pi^{*}(\cdot|s)\|\pi_{\theta_{k}}(\cdot|s))-\mathrm{KL}(\pi^{*}(\cdot|s)\|\pi_{\theta_{k+1}}(\cdot|s))]\right]\end{split}$ (38)
@@ Appendix C Properties of the MLMC Estimates
This section provides some guarantees on the error of the MLMC estimator. This is similar to the results available in (Dorfman & Levy, 2022; Suttle et al., 2023; Beznosikov et al., 2023), although (Dorfman & Levy, 2022; Beznosikov et al., 2023) consider the case of unbiased estimates while our results deal with biased estimates.
@@
Lemma 6 (Lemma 1, (Beznosikov et al., 2023)).
@@ Appendix F Proof of Theorem 3
Lemma 14, (Wei et al., 2020) For any ergodic MDP with mixing time $t_{\mathrm{mix}}$, the following holds for any policy $\pi$.
@@ Appendix H Proof of Theorem 1
$\begin{split}J^{*}-\frac{1}{K}\sum_{k=0}^{K-1}&\operatorname{\mathbb{E}}[J(\theta_{k})]\leq\sqrt{\epsilon_{\mathrm{bias}}}+\frac{G_{1}}{K}\sum_{k=0}^{K-1}\operatorname{\mathbb{E}}\|(\operatorname{\mathbb{E}}_{k}\left[\omega_{k}\right]-\omega^{*}_{k})\|+\dfrac{\alpha G_{2}}{K}\sum_{k=0}^{K-1}\operatorname{\mathbb{E}}\|\omega_{k}-\omega_{k}^{*}\|^{2}\\ &+\dfrac{\alpha\mu^{-2}}{K}\sum_{k=0}^{K-1}\operatorname{\mathbb{E}}\|\nabla_{\theta}J(\theta_{k})\|^{2}+\frac{1}{\alpha K}\operatorname{\mathbb{E}}_{s\sim d^{\pi^{*}}}[\mathrm{KL}(\pi^{*}(\cdot|s)\|\pi_{\theta_{0}}(\cdot|s))].\end{split}$ (58)
@@
We get the desired result by substituting the values of $H,K$ as stated in the theorem. We want to emphasize that the $G_{1}^{2}$ factor with the $\sqrt{\epsilon_{\mathrm{app}}}$ term is a standard component in actor-critic results with a linear critic (Suttle et al., 2023; Patel et al., 2024). However, this factor is often not explicitly mentioned in previous works, whereas we have included it here.